### v1

@@ 3.3 Even-hole free graphs
Figure 5: Proof of Theorem 3. (a) The graph $B_{4}$. (b) The graph $G_{4}$.
It is enough to prove Theorem 3 for each integer $k\geq 4$. For an integer $k\geq 4$, let $B_{k}$ be the graph obtained by taking a triangle $T$ and attaching to one of its vertices a path of length $k$. Let $F_{1}$ denote a copy of $B_{k}$. Let the three vertices of the only triangle of $F_{1}$ be denoted as $a_{1},b_{1},c_{1}$ where the degree of $a_{1}$ is 3, and the only vertex of degree 1 is denoted as $r_{1}$. Let $G_{1}=F_{1}$. For $i=2,\ldots,k$ we construct $G_{i}$ as follows.
  • Let $F_{i}$ denote a copy of $B_{k}$. Let the three vertices of the only triangle of $F_{i}$ are denoted as $a_{i},b_{i},c_{i}$ where the degree of $a_{i}$ is three and the only vertex of degree one is denoted as $r_{i}$.
  • Connect $r_{i}$ and $r_{i-1}$ using a path $P_{i}$ of length $k$. Connect $c_{i}$ and $b_{i-1}$ using a path $Q_{i}$ of length $k+1$. The resulting graph is $G_{i}$.
@@
Let $C$ be an induced cycle of length at least $4$. It is clear from the construction that $C$ will contain at least one vertex from $R$. Moreover, if $r_{i}\in V(C),r_{t}\in V(C)$, then for any $i<j<t$, $r_{j}\in V(C)$. Let $i$ (resp. $t$) be the minimum (resp. maximum) index such that $r_{i}\in V(C)$ (resp. $r_{t}\in V(C)$). Hence, for each $i<j\leq t$, the path $P_{j}$ is also part of $C$. It is also easy to verify that for each $i<j\leq t$, the path $Q_{j}$ is also part of $C$. Let $X_{i}$ (resp. $X_{t}$) be the path between $r_{i}$ and $b_{i}$ (resp. $r_{t}$ and $c_{t}$). Observe that
@@
Hence the length of $C$ is $(t-i)k+(t-i)(k+1)+2(k+1)+(t-i)-1$ which is always odd. ∎
@@
Let $r=r_{k},G=G_{k}$. Then $ipco\left(\overrightarrow{G_{r}}\right)\geq k$.
@@
First, observe that for each $i\in[k]$, $\mathsf{d}\left(r,b_{i}\right)=\mathsf{d}\left(r,c_{i}\right)$. Therefore, if there is an isometric path in $G$ that contains both $b_{i}$ and $c_{i}$, for some $i\in[k]$, then two distinct $r$-rooted isometric paths are needed to cover $b_{i}$ and $c_{i}$. Moreover, it is clear from the construction that for $i<j\in[k]$, there is no directed path from $b_{j}$ to $b_{i}$ in $\overrightarrow{G_{r}}$. Now consider an isometric path $X$ in $G$, and let $t$ be the number of edges of the form $b_{i}c_{i}\in E(X)$. In other words, $t=|\{i\in[k]\colon\{b_{i},c_{i}\}\subseteq V(X)\}|$. Then the above arguments imply that $|A_{r}\left(X\right)|\geq t$. Now we will be done by showing that there is an isometric path that contains at least $k$ many edges of the form $b_{i}c_{i}$.
Let $Q$ be the induced path between $b_{k}$ and $c_{1}$ induced by the vertices of $\left(\displaystyle\bigcup\limits_{i=2}^{k}V(Q_{i})\right)\cup\left(\displaystyle\bigcup\limits_{i=1}^{k}\{b_{j},c_{j}\}\right)$. The length of $Q$ is $k^{2}+k-1$. We claim that $Q$ is an isometric path between $c_{1}$ and $b_{k}$. To prove this, consider a path $Q^{\prime}$ between $c_{1}$ and $b_{k}$ that is different from $Q$. Notice that $Q^{\prime}$ must contain at least one vertex from $R$ and $i$ be the minimum index such that $r_{i}\in R$. Observe that $\mathsf{d}\left(b_{k},r_{i}\right)\geq(k-i+1)k+1$ and $\mathsf{d}\left(r_{i},c_{1}\right)\geq(i+1)k+1$. Therefore, the length of $Q^{\prime}$ is at least $(k+2)k+2$, which is larger than $Q$. Therefore, $Q$ is isometric and contains all edges in $\{b_{i}c_{i}\colon i\in[k]\}$. ∎
Completion of the proof of Theorem 3: Consider two disjoint copies of the graph $G_{k}$, and identify the vertex $r_{k}$ from each copy. Let $Z_{k}$ be the resulting graph. Let $G=G_{k}$ and $r=r_{k}$. Observe that $Z_{k}$ is even-hole free, due to Lemma 4 and the fact that identifying one vertex from two disjoint even-hole free graph results in an even-hole free graph. Moreover, $ipco\left(Z_{k}\right)\geq ipco\left(\overrightarrow{G_{r}}\right)\geq k$, due to Lemma 5.
@@ 4.1 Graph powers
Let $G$ be a graph with $ipco\left(G\right)\leq k$ and $r\geq 2$ be a fixed integer. Let $H=G^{r}$. We will use the following simple observation.
Observation 1.
@@
Let the vertices of $Q$ be ordered as $v_{1}=v,v_{2},\ldots,v_{\ell}$ as they are encountered while traversing $Q$ from $v$. Observe that for $i\in[r]$, the vertices in the set $P_{i}=\{v_{j}\colon j=i+1+xr,x\in\mathbb{N},xr\leq\ell\}$ induce an isometric path in $H$. Since $v$ and $v_{i}$ with $i\in[2,r]$ are adjacent in $H$, we have that $Q_{i}=P_{i}\cup\{v\}$ is also an isometric path in $H$. Now, the set $\mathcal{Q}=\{Q_{i}\colon i\in[r]\}$ covers all the vertices of $Q$. ∎
Now color the edges of $H$ as follows: an edge $e\in E(H)$ is black if $e\in E(G)$; otherwise $e$ is red. An isometric path $P$ of $H$ is red if all edges of $P$ are red.
Input : A graph $G$ and a path $P$ of $G$.
Output : Vertex-minimal non-isometric subpaths of $P$.
1
2if $P$ is isometric in $G$ then
3      Return $\emptyset$.
4else
5       Let $Q$ be any vertex-minimal non-isometric subpath of $P$;
6      $P_{1}$ and $P_{2}$ be the two paths in the subgraph $E(P)\setminus E(Q)$;
7      $Q_{1}=\textsc{Find-Path}(G,P_{1})$
8      $Q_{2}=\textsc{Find-Path}(G,P_{2})$
9      Return $\{Q\}\cup Q_{1}\cup Q_{2}$;
10
Procedure 2 A procedure Find-Path($G,P$) to find vertex-minimal non-isometric paths.
@@ 9      Return $\{Q\}\cup Q_{1}\cup Q_{2}$;
Assume for contradiction that there is a path $P$ in $H$ that has at least two black edges $u_{1}v_{1}$ and $u_{2}v_{2}$. Let the vertices $u_{1},v_{1},u_{2},v_{2}$ appear in that order while traversing $P$ from one end-vertex to the other. Also, assume that the subpath $P^{\prime}$ of $P$ between $u_{1}$ and $v_{2}$ does not contain any other black edge. Clearly, $u_{2}\neq v_{1}$. (Otherwise, the distance between $u_{1}$ and $v_{2}$ is at most 2 in $G$, which implies that $u_{1},v_{2}$ would be adjacent in $H$, contradicting that $P$ is isometric.) Hence, $P^{\prime}$ has at least one edge, and $P^{\prime}$ is a red isometric path in $H$ with $v_{1}$ and $u_{2}$ as its end-vertices. Let $x_{1}=v_{1},x_{2},\ldots,x_{\ell}=u_{2}$ be the vertices of $P^{\prime}$ as they are encountered while traversing $P^{\prime}$ from $v_{1}$ to $u_{2}$. For each $i\in[\ell-1]$, let $P_{i}$ denote an $(x_{i},x_{i+1})$-isometric path of length at most $k$ in $G$. (Since $x_{i}x_{i+1}$ is a red edge, such a path always exists.) For each $i\in[\ell-1]$, let $y_{i}$ denote the vertex of $P_{i}$ that is adjacent to $x_{i}$. Observe that, for each $i\in[\ell-2]$, $y_{i}$ is adjacent to $y_{i+1}$ in $H$. Moreover, $u_{1}$ is adjacent to $y_{1}$ in $H$ and $y_{\ell-1}$ is adjacent to $v_{2}$ in $H$. Hence, $u_{1},y_{1},y_{2},\ldots,y_{\ell-1},v_{2}$ induces a path of length $\ell$ in $H$. But this contradicts the fact that $u_{1},v_{1}=x_{1},\ldots,x_{\ell}=v_{2},u_{2}$ induces an isometric subpath of $P$ with length $\ell+1$. ∎
@@
The vertices of any red isometric path of $H$ can be covered with $12r-$ 1 many isometric paths of $G$.
@@
[…]
Claim 5.14.
@@
Observe that, from the description of Procedure 2 and the definitions, any two paths in $\mathcal{Y}$ are edge disjoint. Hence, the paths in $\mathcal{Y}$ can be ordered as $Y_{1},Y_{2},\ldots,Y_{p}$ such that for $1\leq I<j\leq p$ all edges of $Y_{i}$ appear before those of $Y_{j}$ while traversing from one end-vertex of $Q$ to the other. Without loss of generality, we assume that one end-vertex of $Y_{1}$ is $u_{1}$. Hence, one end-vertex of $Y_{p}$ must be $u_{\ell+1}$. Let the two end-vertices of $Y_{i}$ be denoted as $a_{i}$ and $b_{i}$. Note that $a_{i+1}=b_{i}$ and without loss of generality we can assume that $a_{1}=u_{1}$, and $b_{p}=u_{\ell+1}$. Note that $\sum\limits_{i=1}^{p}|E(Y_{i})|=|E(Q)|\leq\ell\cdot r$. Using the above notations, we prove the following claim.
Claim 5.15.
@@
Observe that $\left|\overline{\mathcal{Z}}\right|\leq\left|\mathcal{Z}\right|+1$. Hence proving $|\mathcal{Z}|\leq r-1$ is enough to prove 5.15. Assume for contradiction that $|\mathcal{Z}|\geq k$. Let $X_{i}$ denote an $(a_{i},b_{i})$-isometric path in $G$ and let $Q^{\prime}$ be the path between $u_{1}$ and $u_{\ell+1}$ obtained by concatenating $X_{1},X_{2},\ldots,X_{p}$. Observe that if for some $i\in[p]$, $Y_{i}\in\overline{\mathcal{Z}}$, then due to 5.14, $|E(X_{i})|=|E(Y_{i})|$. On the other hand, if for some $i\in[p]$, $Y_{i}\in\mathcal{Z}$, then $|E(X_{i})|\leq|E(Y_{i})|-1$. Hence,
@@
$\leq\displaystyle\sum\limits_{Y_{i}\in\overline{\mathcal{Z}}}|E(Y_{i})|+\displaystyle\sum\limits_{Y_{i}\in\mathcal{Z}}|E(Y_{i})|-k$
$\leq\displaystyle\sum\limits_{i}^{p}|E(Y_{i})|-k$
@@
Hence $Q^{\prime}$ is a path between $u_{1}$ and $u_{\ell+1}$ of length at most $r(\ell-1)$. Hence $\mathsf{d}\left(u_{1},u_{\ell+1}\right)$ is at most $r(\ell-1)$. However, this contradicts 1. ∎
Now for each $Y_{i}\in\overline{\mathcal{Z}}$, define $Y^{\prime}_{i}=Y_{i}$ and for each $Y_{i}\in\mathcal{Z}$, define $Y^{\prime}_{i}=Y_{i}\setminus\{b_{i}\}$. Observe that each path in $\mathcal{Y}^{\prime}=\{Y^{\prime}_{i}\colon Y_{i}\in\mathcal{Z}\}\cup\{Y^{\prime}_{i}\colon Y_{i}\in\overline{\mathcal{Z}}\}$ is an isometric path in $G$. Due to 5.15, $|\mathcal{Y}^{\prime}|\leq 2r-1$. Also, $V(P)\subseteq V(Q)\subseteq\displaystyle\bigcup\limits_{Y\in\mathcal{Y}^{\prime}}V(Y)$. Therefore, the vertices of $P$ can be covered with $12k-$ 1 many isometric paths in $G$. ∎
Completion of the proof of Theorem 4(a): Let $G$ be a graph with $ipco\left(G\right)\leq k$ and $H$ be the $r^{th}$ power of $G$. Let $v$ be a vertex of $G$ such that the vertices of any isometric path of $G$ can be covered with $k$ many $v$-rooted isometric paths. Let $P$ be an isometric path of $H$. Due to Lemma 7 $P$ contains at most one black edge $e$. Let $P_{1}$ and $P_{2}$ be the two connected components obtained by removing $e$ from $P$. Note that both $P_{1}$ and $P_{2}$ are red isometric paths. Due to Lemma 8, for each $i\in\{1,2\}$, the vertices of $P_{i}$ can be covered with $12r-$ 1 many $v$-rooted isometric paths of $G$. Hence, the vertices of $P$ can be covered by $24r-$ 2 $v$-rooted isometric paths of $G$. This implies that the vertices of $P$ can be covered with $k(4r-2)$ many $v$-rooted isometric paths of $G$. Thus, Lemma 6 implies that $P$ can be covered by $k(4r^{2}-2r)$ $v$-rooted isometric paths of $H$. Hence, $ipco\left(H\right)\leq k(4r^{2}-2r)$. Similar arguments as above imply that if $sipco\left(G\right)\leq k$, then $sipco\left(H\right)\leq k{(4r^{2}-2r)}$.

### v2

@@ 3.3 Even-hole free graphs
Figure 9: Proof of Theorem˜7. (a) The graph $B_{4}$. (b) The graph $G_{4}$.
In this section, we show Theorem˜7 that for each integer $k\geq 1$, there exists an even-hole free graph $Z_{k}$ with maximum degree at most $3$, such that $ipco\left(Z_{k}\right)\geq k$. It is enough to prove the statement for each integer $k\geq 4$ (and setting $Z_{i}=Z_{4}$ for $i=1,2,3$). For an integer $k\geq 4$, let $B_{k}$ be the graph obtained by taking a triangle $T$ and attaching to one of its vertices a path of length $k$ (see Figure˜9. We construct $G_{k}$ as follows.
  • Let $F_{1},F_{2},\ldots,F_{k}$ be disjoint copies of $B_{k}$. Let $a_{i},b_{i},c_{i}$ be the three vertices of the unique triangle of $F_{i}$, where the degree of $a_{i}$ is 3, and the only vertex of degree 1 in $F_{i}$ is denoted by $r_{i}$.
  • Connect $r_{i}$ and $r_{i-1}$ using a path $P_{i}$ of length $k$. Connect $c_{i}$ and $b_{i-1}$ using a path $Q_{i}$ of length $k+1$. The resulting graph is $G_{k}$.
@@
Let $C$ be an induced cycle of length at least $4$. It is clear from the construction that $C$ will contain at least one vertex from $R$. Moreover, if $r_{i},r_{t}\in V(C)$ for some $i<t$, then for any $i<j<t$, $r_{j}\in V(C)$. Let $i$ (resp. $t$) be the minimum (resp. maximum) index such that $r_{i}\in V(C)$ (resp. $r_{t}\in V(C)$). Hence, for each $i<j\leq t$, the path $P_{j}$ is also part of $C$. It is also easy to verify that for each $i<j\leq t$, the path $Q_{j}$ is also part of $C$. Let $X_{i}$ (resp. $X_{t}$) be the path between $r_{i}$ and $b_{i}$ (resp. $r_{t}$ and $c_{t}$). Observe that
@@
Hence the length of $C$ is (ti)k+(ti)(k+1)+2(k+1)+(ti)1=(ti+1)(2k+2)1(t-i)k+(t-i)(k+1)+2(k+1)+(t-i)-1=(t-i+1)(2k+2)-1, which is always odd. ∎
@@
Let $r=r_{k},G=G_{k}$. Then $ipco\left(G,r\right)\geq k$.
@@
Let $Q$ be the induced path between $b_{k}$ and $c_{1}$ induced by the vertices of $\left(\displaystyle\bigcup\limits_{i=2}^{k}V(Q_{i})\right)\cup\left(\displaystyle\bigcup\limits_{i=1}^{k}\{b_{j},c_{j}\}\right)$. The length of $Q$ is $k^{2}+k-1$. We claim that $Q$ is an isometric path between $c_{1}$ and $b_{k}$. To prove this, consider a path $Q^{\prime}$ between $c_{1}$ and $b_{k}$ that is different from $Q$. Notice that $Q^{\prime}$ must contain at least one vertex from $R$ and let $i$ be the minimum index such that $r_{i}\in R$. Observe that $\mathsf{d}\left(b_{k},r_{i}\right)\geq(k-i+1)k+1$ and $\mathsf{d}\left(r_{i},c_{1}\right)\geq(i+1)k+1$. Therefore, the length of $Q^{\prime}$ is at least $(k+2)k+2$, which is larger than the length of $Q$. Therefore, $Q$ is isometric and contains all edges in $\{b_{i}c_{i}\colon i\in[k]\}$.
Now, observe that for each $i\in[k]$, $\mathsf{d}\left(r,b_{i}\right)=\mathsf{d}\left(r,c_{i}\right)$. Therefore, two distinct $r$-rooted isometric paths are needed to cover $b_{i}$ and $c_{i}$, for $i\in[k]$. Moreover, it is clear from the construction that for $i<j\in[k]$, there is no directed path from $b_{j}$ to $b_{i}$ nor from $b_{i}$ to $b_{j}$ in $\overrightarrow{G_{r}}$. The above arguments imply that $|A_{r}\left(Q\right)|\geq k$ and therefore $ipco\left(G,r\right)\geq k$. ∎
Completion of the proof of Theorem˜7. Consider two disjoint copies $X,Y$ of the graph $G_{k}$. Let $r_{1}$ and $r_{2}$ denote the vertices labeled $r_{k}$ in $X$ and $Y$, respectively. Now, add an edge between the vertices $r_{1}$ and $r_{2}$. Let $Z_{k}$ be the resulting graph. Observe that $Z_{k}$ is even-hole free, due to Lemma˜16 and the fact that adding an edge between two disjoint even-hole free graphs results in an even-hole free graph. Let $v$ be any vertex of $Z_{k}$ and without loss of generality, $v\in V(X)$. Observe that, for any vertex $u\in V(Y)$, the $(u,v)$-isometric path contains the vertex $r_{2}$. Hence, for any isometric path $P$ that lies in $Y$, a set of $v$-rooted isometric paths that cover $P$, is also a set of $r_{2}$-rooted isometric paths. Hence, $ipco\left(Z_{k},v\right)\geq ipco\left(Y,r_{2}\right)\geq k$.
Remark 1.
The graphs constructed to prove Theorem˜7 can also be used to show that the property of bounded (strong) isometric path complexity is not preserved under quasi-isometry. Indeed, consider the graph $G_{k}$ and the $(c_{1},b_{k})$-isometric path that does not go through $r_{1}$. By appropriately subdividing some edges of this path, it is possible to obtain a graph $H$ which is quasi-isometric to $G$ but has constant (strong) isometric path complexity.
@@ 4.1 Graph powers
Let $G$ be a graph with $ipco\left(G\right)\leq k$ and $r\geq 2$ be a fixed integer. Let $H=G^{r}$. Now color the edges of $H$ as follows: an edge $e\in E(H)$ is black if $e\in E(G)$; otherwise $e$ is red. An isometric path $P$ of $H$ is red if all edges of $P$ are red. We will use the following simple observation.
Observation 18.
@@
Let the vertices of $Q$ be ordered as $v_{1}=v,v_{2},\ldots,v_{\ell}$ as they are encountered while traversing $Q$ from $v$. Observe that for $i\in[r]$, the vertices in the set $P_{i}=\{v_{j}\colon j=i+1+xr,x\in\mathbb{N},i+1+xr\leq\ell\}$ induce an isometric path in $H$. Since $v$ and $v_{i}$ with $i\in[2,r]$ are adjacent in $H$, we have that $Q_{i}=P_{i}\cup\{v\}$ is also an isometric path in $H$. Now, the set $\mathcal{Q}=\{Q_{i}\colon i\in[r]\}$ covers all the vertices of $Q$. ∎
@@
Assume for contradiction that there is a path $P$ in $H$ that has at least two black edges $u_{1}v_{1}$ and $u_{2}v_{2}$. Let the vertices $u_{1},v_{1},u_{2},v_{2}$ appear in that order while traversing $P$ from one end-vertex to the other. Also, assume that the subpath $P^{\prime}$ of $P$ between $u_{1}$ and $v_{2}$ does not contain any other black edge. Clearly, $u_{2}\neq v_{1}$. (Otherwise, the distance between $u_{1}$ and $v_{2}$ is at most 2 in $G$, which implies that $u_{1},v_{2}$ would be adjacent in $H$, contradicting that $P$ is isometric.) Hence, $P^{\prime}$ has at least one edge, and $P^{\prime}$ is a red isometric path in $H$ with $v_{1}$ and $u_{2}$ as its end-vertices. Let $x_{1}=v_{1},x_{2},\ldots,x_{\ell}=u_{2}$ be the vertices of $P^{\prime}$ as they are encountered while traversing $P^{\prime}$ from $v_{1}$ to $u_{2}$. For each $i\in[\ell-1]$, let $P_{i}$ denote an $(x_{i},x_{i+1})$-isometric path of length at most $r$ in $G$. (Since $x_{i}x_{i+1}$ is a red edge, such a path always exists.) For each $i\in[\ell-1]$, let $y_{i}$ denote the vertex of $P_{i}$ that is adjacent to $x_{i}$. Observe that, for each $i\in[\ell-2]$, $y_{i}$ is adjacent to $y_{i+1}$ in $H$. Moreover, $u_{1}$ is adjacent to $y_{1}$ in $H$ and $y_{\ell-1}$ is adjacent to $v_{2}$ in $H$. Hence, $u_{1},y_{1},y_{2},\ldots,y_{\ell-1},v_{2}$ induces a path of length $\ell$ in $H$. But this contradicts the fact that $u_{1},v_{1}=x_{1},\ldots,x_{\ell}=v_{2},u_{2}$ induces an isometric subpath of $P$ with length $\ell+1$. ∎
@@
The vertices of any red isometric path of $H$ can be covered by 2r12r-1 many isometric paths of $G$.
@@
Let $\mathcal{Z}$ be a maximal collection of edge-disjoint non-isometric subpaths of $Q$, and $\overline{\mathcal{Z}}$ is the set of connected components in the graph obtained by removing all paths in $\mathcal{Z}$ from $Q$. Let $\mathcal{Y}=\mathcal{Z}\cup\overline{\mathcal{Z}}$. We shall also use the following.
Claim 4.1.N.
@@
The definitions imply that any two paths in $\mathcal{Y}$ are edge disjoint. Hence, the paths in $\mathcal{Y}$ can be ordered as $Y_{1},Y_{2},\ldots,Y_{p}$ such that for $1\leq i<j\leq p$ all edges of $Y_{i}$ appear before those of $Y_{j}$ while traversing from one end-vertex of $Q$ to the other. Hence, one end-vertex of $Y_{p}$ must be $u_{\ell+1}$. Let the two end-vertices of $Y_{i}$ be denoted as $a_{i}$ and $b_{i}$. Note that $a_{i+1}=b_{i}$ and without loss of generality we can assume that $a_{1}=u_{1}$, and $b_{p}=u_{\ell+1}$. Note that $\sum\limits_{i=1}^{p}|E(Y_{i})|=|E(Q)|\leq\ell\cdot r$. Using the above notations, we prove the following claim.
Claim 4.1.O.
@@
Observe that $\left|\overline{\mathcal{Z}}\right|\leq\left|\mathcal{Z}\right|+1$. Hence proving $|\mathcal{Z}|\leq r-1$ is enough to prove ˜4.1.O. Assume for contradiction that $|\mathcal{Z}|\geq r$. Let $X_{i}$ denote an $(a_{i},b_{i})$-isometric path in $G$ and let $Q^{\prime}$ be the path between $u_{1}$ and $u_{\ell+1}$ obtained by concatenating $X_{1},X_{2},\ldots,X_{p}$. Observe that for each $i\in[p]$, $Y_{i}\in\overline{\mathcal{Z}}$, then due to ˜4.1.N, $|E(X_{i})|=|E(Y_{i})|$. On the other hand, if for each $i\in[p]$, $Y_{i}\in\mathcal{Z}$, then $|E(X_{i})|\leq|E(Y_{i})|-1$. Hence,
@@
$\leq\displaystyle\sum\limits_{Y_{i}\in\overline{\mathcal{Z}}}|E(Y_{i})|+\displaystyle\sum\limits_{Y_{i}\in\mathcal{Z}}|E(Y_{i})|-r$
$\leq\displaystyle\sum\limits_{i=1}^{p}|E(Y_{i})|-r$
@@
Hence $Q^{\prime}$ is a path between $u_{1}$ and $u_{\ell+1}$ of length at most $r(\ell-1)$. Hence $\mathsf{d}\left(u_{1},u_{\ell+1}\right)$ is at most $r(\ell-1)$. However, this contradicts ˜18. ∎
Now for each $Y_{i}\in\overline{\mathcal{Z}}$, define $Y^{\prime}_{i}=Y_{i}$ and for each $Y_{i}\in\mathcal{Z}$, define $Y^{\prime}_{i}=Y_{i}\setminus\{b_{i}\}$. Observe that each path in $\mathcal{Y}^{\prime}=\{Y^{\prime}_{i}\colon Y_{i}\in\mathcal{Z}\}\cup\{Y^{\prime}_{i}\colon Y_{i}\in\overline{\mathcal{Z}}\}$ is an isometric path in $G$. Due to ˜4.1.O, $|\mathcal{Y}^{\prime}|\leq 2r-1$. Also, $V(P)\subseteq V(Q)\subseteq\displaystyle\bigcup\limits_{Y\in\mathcal{Y}^{\prime}}V(Y)$. Therefore, the vertices of $P$ can be covered by 2r12r-1 many isometric paths in $G$. ∎
Completion of the proof of Theorem˜8(a). Let $G$ be a graph and $H$ be the $r^{th}$ power of $G$. Let $v$ be a vertex of $G$, and $k=ipco\left(G,v\right)$. Let $P$ be an isometric path of $H$. Due to Lemma˜20, $P$ contains at most one black edge $e$. Let $P_{1}$ and $P_{2}$ be the two connected components obtained by removing $e$ from $P$. Note that both $P_{1}$ and $P_{2}$ are red isometric paths. Due to Lemma˜21, for each $i\in\{1,2\}$, the vertices of $P_{i}$ can be covered by 2r12r-1 isometric paths of $G$. Hence, the vertices of $P$ can be covered by 4r24r-2 isometric paths of $G$. This implies that the vertices of $P$ can be covered by k(4r2)k(4r-2) many $v$-rooted isometric paths of $G$. Thus, Lemma˜19 implies that $P$ can be covered by $k(4r^{2}-2r)$ $v$-rooted isometric paths of $H$. The above arguments imply $sipco\left(H\right)\leq(4r^{2}-2r)sipco\left(G\right)$ and $ipco\left(H\right)\leq(4r^{2}-2r)ipco\left(G\right)$.